OpenStudy (anonymous):

What is the limit of: root (2x+1) minus root 3 divided by (x-1)

zepdrix (zepdrix):

Limit as x approaches `what value` ? :)

OpenStudy (anonymous):

as x approaches 1 from the right

zepdrix (zepdrix):

\[\Large \lim_{x\to1^+}\frac{\sqrt{2x+1}-\sqrt3}{x-1}\]

zepdrix (zepdrix):

We want to start by multiplying the top and bottom by the `conjugate` of the numerator,\[\Large \color{orangered}{\sqrt{2x+1}+\sqrt3}\]k?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Leave the mess alone in the denominator. We should be able to clean up the top though. So multiplying conjugates gives us the `difference of squares` right? remember that? :o

zepdrix (zepdrix):

Here is how that works: \[\large (a-b)(a+b) \quad=\quad a^2-b^2\] But for our problem here, it'll look more like this:\[\large (\sqrt a-\sqrt b)(\sqrt a+\sqrt b) \quad=\quad a-b\]

zepdrix (zepdrix):

\[\Large \lim_{x\to1^+}\frac{\sqrt{2x+1}-\sqrt3}{x-1}\color{orangered}{\left(\frac{\sqrt{2x+1}-\sqrt3}{\sqrt{2x+1}-\sqrt3}\right)}\] Understand how this numerator will multiply out?

zepdrix (zepdrix):

Grrr typo... sec

OpenStudy (anonymous):

No worries about the typo haha. Will the numerator end up being (2x+1) minus root 9?

zepdrix (zepdrix):

yah looks good.

zepdrix (zepdrix):

Simplify it down further! :)

OpenStudy (anonymous):

simplified to 2(x-1) ?

zepdrix (zepdrix):

niceeee

zepdrix (zepdrix):

\[\Large \lim_{x\to1^+}\frac{2(x+1)}{(x+1)(stuff)}\] Looks like you're on the right track :) See a cancellation from there?

zepdrix (zepdrix):

Omg I can't type correctly today lol

zepdrix (zepdrix):

\[\Large \lim_{x\to1^+}\frac{2(x-1)}{(x-1)(stuff)}\]

OpenStudy (anonymous):

the two (x-1)s will cancel out leaving 2 over all the stuff?

zepdrix (zepdrix):

Ya! And since we done some `cancelling`, we should now check to see if 1 is still giving us a problem.

zepdrix (zepdrix):

At the start of the problem, if we tried to plug x=1 directly in, we ended up with the indeterminate form 0/0. <- bad. What happens if we plug x=1 directly in from where we are now?

OpenStudy (anonymous):

I ended up getting 2/ root 5 + root 3 but I don't think that's right

zepdrix (zepdrix):

Hmm, yah I think you made a boo boo in there somewhere.\[\Large \frac{2}{\sqrt{2(1)+1}+\sqrt3}\quad = \quad \frac{2}{\sqrt3+\sqrt3}\]

OpenStudy (anonymous):

OH I plugged 2 in instead of 1! Okay now I got 2 over root 3 plus root 3. There has to be something further I'm guessing because the book says the answer is root 3 over 3

zepdrix (zepdrix):

So adding the terms in the denominator:\[\Large \frac{2}{2\sqrt3} \quad=\quad \frac{1}{\sqrt3}\]Right?

zepdrix (zepdrix):

It's usually a good idea to try and get `irrational numbers` out of the denominator when possible. So, from where we are, we want to move the sqrt3 up to the numerator with multiplication.\[\Large \frac{1}{\sqrt3}\cdot\frac{\sqrt3}{\sqrt3} \quad=\quad ?\]

OpenStudy (anonymous):

wait, how did you get 2 over 2root3?

zepdrix (zepdrix):

2 apples, right? :o when we add them